# Practical implementation of a single-qubit rotation algorithm

### v1

@@ Abstract
The Toffoli is an important universal quantum gate, and will alongside the Clifford gates be available in future Fault-Tolerant Quantum Computing hardware. Many quantum algorithms rely on performing arbitrarily small single-qubit rotations for their function, and these rotations may also be used to construct any unitary from a limited (but universal) gate set; it is then of significant interest how to carry out such rotations. In this work, we evaluate the performance of a recently proposed single-qubit rotation algorithm using the Clifford+Toffoli gate set by implementation on both a real and simulated quantum computer. We test the algorithm under various simulated noise levels utilizing a per-qubit depolarizing error noise model, finding that the errors are seemingly explained by a binomial distribution wherein these errors change controlling ancilla measurements from 00 into $1$. Similar observations appear to hold when conducting live runs; noise levels here make further meaningful conclusions difficult, although for the smallest possible number of ancillary controls we do note that error mitigation is helpful. Our results suggest that the algorithm will perform well under up to $1\%$ noise, under the noise model we chose. Our results also suggest the algorithm could be used as a benchmark for Quantum Processing Units, given its linear increase in total number of qubits and Toffoli gates required.
@@ I Introduction
The Clifford+Toffoli gate set will be available in future quantum computer hardware: It is therefore natural to ask if generic unitaries can be built using that gate set, and indeed, if they can be built efficiently and explicitly. We were recently able to answer this question in the affirmative, by giving an explicit construction for approximating any $z$-axis rotation to within distance $\epsilon$ using the Clifford+Toffoli gate set with expected Toffoli count $4\lceil\log\frac{1}{\epsilon}\rceil$, see Ref. [11]: Three such rotations allows approximation of any single qubit gate. Our construction is simple and efficient, requiring only a trigonometric expression of the desired rotation angle, rounded to the desired accuracy. The resulting algorithm succeeds in applying a rotation $\epsilon$-close to that desired with probability strictly greater than $1/2$, has expected gate depth $34\lceil\log\frac{1}{\epsilon}\rceil+$ 3 and uses $2\lceil\log\frac{1}{\epsilon}\rceil$ ancillas. Should the rotation not be applied, a $Z$ gate is applied instead, and ancilla measurements indicate this.
@@ II The single-qubit rotation algorithm
We briefly describe the algorithm implemented in this work; for more details, we refer to Ref. [11]. Let $R_{\theta}$ denote a single-qubit rotation (along the $z$-axis) of angle $\theta$ radians. A simple example taken from [13, p. 198], of a circuit implementing the rotation $R_{\arccos 3/5}$, can be found in Figure 1a. This circuit can then be generalized as depicted in Figure 1b. Another way of looking at the circuit in Figure 1a is that it performs a comparison with the classical constant $k=3$. Increasing the number of ancillary controls $n$ then allows us to set other $k$: Say that we wish to apply a specific rotation of angle $\theta$ to within distance $\epsilon$. Then by Theorem 1 of Ref. [11], we should choose a number of ancillary controls equal to the (smallest) integer $n$ such that
@@ V.1 Simulated runs
Figure 6: Probability of applying $T^{*}$, live runs.
Figure 7: PF calculated for $T^{*}$ and $T$, live runs.
Figure 8: PF calculated for $Z^{*}$ and $Z$, live runs.
@@ VI Conclusions
Our algorithm does work under modest noise levels, though such (low) noise levels that are required seem not presently available: This is unfortunate but expected, as our algorithm is primarily intended for use on FTQC devices, rather than Noisy Intermediate Scale Quantum (NISQ) computational devices. Hopefully, our conjectured means of resolving this as stated above may improve our algorithm’s performance on NISQ devices somewhat: It seems to us to at least serve a useful purpose as a benchmark, offering almost step-by-step scaling of qubits and a fairly straightforward means of verifying the output using quantum tomography.

### v2

@@ Abstract
The Toffoli is an important universal quantum gate, and will alongside the Clifford gates be available in future Fault-Tolerant Quantum Computing hardware. Many quantum algorithms rely on performing arbitrarily small single-qubit rotations for their function, and these rotations may also be used to construct any unitary from a limited (but universal) gate set; it is then of significant interest how to carry out such rotations. In this work, we evaluate the performance of a recently proposed single-qubit rotation algorithm using the Clifford+Toffoli gate set by implementation of a one-shot version on both a real and simulated quantum computer. We test the algorithm under various simulated noise levels utilizing a per-qubit depolarizing error noise model, finding that the errors are seemingly explained by a binomial distribution wherein these errors change controlling ancilla measurements from 00 into $1$. Similar observations appear to hold when conducting live runs; noise levels here make further meaningful conclusions difficult, although for the smallest possible number of ancillary controls $n$ we do note that error mitigation is helpful. Our results suggest that the algorithm will perform well under up to $1\%$ noise, under the noise model we chose. We further posit the use of our algorithm as a benchmark for Quantum Processing Units, given that its one-shot version has a very modest (in fact close to minimal) linear increase in total number of qubits and Toffoli gates required when increasing $n$: We provide details on how to go about this.
@@ I Introduction
The Clifford+Toffoli gate set will be available in future quantum computer hardware: It is therefore natural to ask if generic unitaries can be built using that gate set, and indeed, if they can be built efficiently and explicitly. We were recently able to answer this question in the affirmative, by giving an explicit construction for approximating any $z$-axis rotation to within distance $\epsilon$ using the Clifford+Toffoli gate set with expected Toffoli count $4\lceil\log\frac{1}{\epsilon}\rceil$, see Ref. [11]: Three such rotations allows approximation of any single qubit gate. Our construction is simple and efficient, requiring only a trigonometric expression of the desired rotation angle, rounded to the desired accuracy. The resulting algorithm succeeds in applying a rotation $\epsilon$-close to that desired with probability strictly greater than $1/2$, has expected gate depth $64\lceil\log\frac{1}{\epsilon}\rceil+$ 6 and uses $2\lceil\log\frac{1}{\epsilon}\rceil$ ancillas. Should the rotation not be applied, a $Z$ gate is applied instead, and ancilla measurements indicate this.
@@
Beyond its usefulness in an FTQC device, our algorithm also demonstrates a linear scaling in its requisite number of ancilla and Toffoli gates. This paves the way for it to be used as a remarkably simple means of benchmarking for comparing Noisy Intermediate Scale Quantum (NISQ) devices: The number of qubits in active use along with number of universal gates are typically considered the greatest contributors of noise within such a device [10]. Typically our algorithm is utilized by first selecting a tolerable error rate $\epsilon$ which decides the number of ancillary controls $n$; here one then chooses the smallest such $n$. There is, however, nothing preventing a reversed means of application: Select first an $n$ disregarding error considerations, which then dictates the $\epsilon$. In so doing, the total number of qubits in active use by the algorithm will be set to $22n-$ 2, plus a single qubit as target for the rotation, and the total number of required Toffolis will also be $22n-$ 2. In other words, increasing $n$ by $1$ increases total number of qubits and Toffolis by $2$ each. The actual benchmarking can then consist of observing the actual probability of the rotation being applied as compared to its theoretical value, or reconstructing the noisy channel via quantum state [13] and process [14] tomography and comparing it against a suitable unitary, or both of these in tandem: We provide an example of how to go about the latter by explicit implementation in this work.
@@ II The single-qubit rotation algorithm
We briefly describe the algorithm implemented in this work; for more details, we refer to Ref. [11]. Let $R_{\theta}$ denote a single-qubit rotation (along the $z$-axis) of angle $\theta$ radians. A simple example taken from [15, p. 198], of a circuit implementing the rotation $R_{\arccos 3/5}$, can be found in Figure 1a. This circuit can then be generalized as depicted in Figure 1b. Another way of looking at the circuit in Figure 1a is that it performs a comparison with the classical constant $k=3$. Increasing the number of ancillary controls $n$ then allows us to set other $k$: Say that we wish to apply a specific rotation of angle $\theta$ to within distance $\epsilon$. Then by Theorem 1 of Ref. [11], we should choose a number of ancillary controls equal to the (smallest) integer $n$ such that
@@ V Results
Figure 6: Probability of applying $T^{*}$, live runs.
Figure 7: PF calculated for $T^{*}$ and $T$, live runs.
Figure 8: PF calculated for $Z^{*}$ and $Z$, live runs.
@@ VI Conclusions
Our algorithm does work under modest noise levels, though such (low) noise levels that are required seem not presently available: This is unfortunate but expected, as our algorithm is primarily intended for use on FTQC devices, rather than Noisy Intermediate Scale Quantum (NISQ) computational devices. Hopefully, our conjectured means of resolving this as stated above may improve our algorithm’s performance on NISQ devices somewhat. This seems to us a good starting point if utilizing our algorithm as a benchmark: It may well be that some relatively simple additional error-mitigation techniques will allow for good results when $n=4$, at which point $n=5$ is the next step, etc. We would also note that even though technically the algorithm does reduce in complexity whenever $k$ is even (recall the discussion on circuit simplifications previously), such a reduction is not mandated: We can construct the explicit circuit for any $n$, should we so desire. This might then offer a tantalizing alternative to going straight from $n=2$ to $n=4$ in our example: Construct and run the circuit for $n=3$ instead, and see if this indeed yields more favorable results compared to $n=4$, by the means described above, i.e. simple comparisons of the probability of applying a rotation in conjunction with quantum tomography.